OpenStudy (anonymous):

Solve the following differential equations (show your solutions): a. dx/dt=r , x(0)=a b. dx/dt=rx , x(0)=a c. dx/dt=rx(1-x) , x(0)=a

OpenStudy (blockcolder):

All of them are separable. For the third one, you require partial fractions.

OpenStudy (anonymous):

solution please?

OpenStudy (blockcolder):

\[\frac{dx}{dt}=r\\ dx=r\ dt\\ x=rt+C\\ a=r(0)+C=C \Rightarrow x=rt+a\]

OpenStudy (anonymous):

how about letter c?

OpenStudy (blockcolder):

\[\frac{dx}{dt}=rx\\ \frac{dx}{x}=r\ dt\\ \ln|x|=rt+C\\ \ln|a|=r(0)+C=C\Rightarrow \ln|x|=rt+\ln|a|\\ \ln|x|-\ln|a|=rt\\ \left |\frac{x}{a}\right |=e^{rt}\\ |x|=|a|e^{rt}\]